OpenStudy (anonymous):

Solution A is 30% antifreeze and Solution B is 60% antifreeze. How many gallons of each Solution B must be combined with Solution A in order to produce an 18 gallon mixture that iis 50% antifreeze?

OpenStudy (anonymous):

A+B= 18 A=18 -B (A*0.3)+(B*0.6) = (18*0.5) Subtitution : A=18 -B ((18 -B)*0.3)+(B*0.6) = (18*0.5) can you solve from now?